Left Direct Gregorian Technical Observing with IRTC: UT 2009 October 09

Observer: JHill (Tucson), DMiller (LBTO)
Telescope Operator: Aaron Ceranski
SW support: TLeibold (Tucson)

Summary

A very successful nignt. We fixed an x-flip problem in GCS, a pcsmx flip in oacontrol and a wfsc_pupil_shift_x flip in GCS

Collected center of rotation data, WFSC pinhole position data, a small pointing model data set, and images of the moon (with no appearance of a plume from the LCROSS)

The seeing ranged from 1" to 2" all night, with the exception of 0.7-0.8" for the last hour.

Details

All Times in UT

00:57 presetTelescope BS9185 Successful up to GCS not finding a guide star

01:05 Aaron Opens up chamber and points East. The primary is at 7 C and ambient is -1 C. Dome seeing will be bad for awhile.

01:06 several single images with IRTC in /Repository/AGW_Data/091090iraf

01:14 cube of 100 "almost" sky flats. It's pretty dark now

01:16 presetTelescope M5_1539 ACQUIRE to set IE and CA. No star seen

01:25 ptspiral Star found

01:28 set IE=-16 CA=37

01:31 presetTelescope BS9177 ACQUIRE Star on edge. Set IE=-35 CA=47.

Guiding

Seeing ~2.0"

01:40 presetTelescope BS9177 ACTIVE gs=0 Star acquired, but pointing correction did not move the correct direction. Guide star was not in the guide box. Torsten changed pointingCorr_flip_x and guidingCorr_flip_x both to +1 in AIP_IRTC_DG_L.cfg. New prest will read this file

01:50 presetTelescope BS9177 ACTIVE gs=0 Star positioned and guiding. +1,+1 is what we needed. Letting WFSC collimate. WFSC images are about as wavey as I have seen.

02:00 presetTelescope BS9177 ACTIVE gs=41 Star acquired and guiding started. On the acquisition image, the star was 12" to the right of the hotspot. Probably pointing drift from the last IE CA setting

2:07 presetTelescope BS9177 ACQUIRE gs=0 Star was 12" to the left. Set IE=-43 CA=53

2:14 presetTelescope BS9177 ACTIVE gs=41 PA=168 Star acquired and guiding started. Star was 2" to the right. This is surely point drift.

2:15 Humidity dropped from 70% to 10% in the last 30 minutes

2:20 presetTelescope BS9177 ACTIVE gs=41 PA=120 No star in acquisition image. Torsten change pcsmx from -1 to +1

2:30 presetTelescope BS9177 ACQUIRE gs=0 Star in the middle IE CA not changed

2:32 presetTelescope BS9177 ACTIVE gs=41 PA=120 No Star

2:32 presetTelescope BS9177 ACTIVE gs=41 PA=168.77 Star found down and to the right by 2"

2:35 presetTelescope BS9177 ACTIVE gs=0 PA=168.77 Star found up and to the right by 1"

2:42 presetTelescope BS9177 ACTIVE gs=41 PA=168.77 Star found to the right by 2"

2:43 presetTelescope BS9177 ACTIVE gs=41 PA=166.77 PCS_XY = -4.0,129.9 Star was 15" to the left on the acquisition image.

2:52 Torsten dhange pcsmx back to -1 restart aocontrol

2:56 presetTelescope BS9177 ACQUIRE gs=0 PA=168.77 out of focus

2:58 presetTelescope BS9177 ACTIVE gs=41 PA=167.25 Star moved the same diretion (left) on the acquisition image. Thus, it looks like oaccontrol is not using the pcsmx parameter

3:10 John finds in the log that GCS is using it's internal transformation rather that the one calculated by oaccontrol. Torsten realized that the problem with OSS and the fix of using an old library is the cause. The bug in the new library causes an error when a boolean parameter is read from file. If there is an error, then the default value is used, which is for GCS is to use it's internal transformation.

3:15 presetTelescope BS9177 ACTIVE gs=41 PA=167.25 Star moved the same direction (left). However, John found in the log that GCS is still using it's internal transformation.

3:23 Torsten is going to ship a new build of GCS to the mountain with the external transformation as the default

3:34 presetTelescope BS9177 ACTIVE gs=41 PA=167.25 Star moved the same direction (left). However, GCS is now using the external transformation from aoccontrol.

3:35 Torsten change pcsmx to +1 and restart. Oops, sign was not change. Torsten try again.

3:48 presetTelescope BS9177 ACTIVE gs=41 PA=167.25 Star found a ~4" down (this is Y direction and probably drift)

3:50 presetTelescope BS9177 ACTIVE gs=0 PA=167.25 Star found a ~4" down, same as off-axis star above

3:52 presetTelescope BS9177 ACQUIRE gs=0 PA=167.25 IE=-41 CA=56

3:54 presetTelescope BS9177 ACTIVE gs=41 PA=167.25 Star is within 1" and is down (direction of drift)

3:54 presetTelescope BS9177 ACTIVE gs=41 PA=120 Star is within 1".

X-Flips are now correct.

Active Optics

Seeing ~1.1"

4:15 After collimating for awhile, it is clear that there is a problem with Active Optics off the y-axis. May be the due to non-zero x is wrong. I have a look with my IDL routines, and the pupils is in the wrong place. Check with an on the y-axis star

4:20 presetTelescope BS9177 ACTIVE gs=45 Find that the .info file that Torsen is writing has different values than those written to the log. The log values are close to my IDL values.

4:20 presetTelescope BS9177 ACTIVE gs=45 PA=10 (45 from the PA=55 above) Pupil position is about the same. Converges to 250 nm RMS

4:20 presetTelescope BS9177 ACTIVE gs=45 PA=-20 The IDL and GCS pupil_x positions were different by a subap. John set WFS_pupil_shift_x to +1 from -1

4:20 presetTelescope BS9177 ACTIVE gs=45 PA=-20 Now IDL and GCS are within a fraction of a pixel. NOTE Torsten writes the on-axis pupil_x and pupil_y values to the .info file SHOULD BE FIXED

4:20 presetTelescope BS9177 ACTIVE gs=45 PA=-20 Collimate for awhile. Z10 seems to John to be oscillating. Aaron put +500 Z10 on the primary. This showed up only in Z10 and in two iterations -450 had been removed.

4:50 presetTelescope BS9177 ACTIVE gs=41 PA=167.25 Collimation looks good. Small corrections.

Active Optics as working well

Center of rotation of Guide Probe

Seeing ~1.5"

5:10 presetTelescope BS9101 ACTIVE gs=0 No star found

5:16 presetTelescope M5_0036 ACQUIRE gs=0 IE=-42 CA=40

5:20 presetTelescope BS9101 ACTIVE gs=0 Collimate

5:40 Turn off guiding and collimation

5:43 Take 70 sec guider image with rotator going from -43 to 320 /Repository/AGW_Data/091009guider/guider000007.fits

5:55 320 to -40 /Repository/AGW_Data/091009guider/guider000008.fits

6:00 -40 to 320 /Repository/AGW_Data/091009guider/guider000009.fits

Center of rotation of IRTC

Seeing ~1.2"

6:08 presetTelescope M5_0036 ACQUIRE gs=5 Oops, wanted ACTIVE

6:09 presetTelescope M5_0036 ACTIVE gs=5 Collimate

6:17 rotate from 60 to 440 with 10ms IRTC image at 10 Hz (800 Frames) /Repository/IRTC_Data/default/irtc_00002_cube.fit

  • Results for irtc_00002_cube.fit
    • Center (X,Y) = (162.75,131.5)
    • Center of Rotation of AGw#4 (irtc_00002_stacked.fits):
      irtc_00002_stacked.jpg

6:20 rotate from 440 to 60 (1000 Frames) /Repository/IRTC_Data/default/irtc_00003_cube.fit

  • Results for irtc_00003_cube.fit
    • Center (X,Y) = (162.0,131.75)
    • Center of Rotation of AGw#4 (irtc_00003_stacked.fits):
      irtc_00003_stacked.jpg

6:23 rotate from 440 to 60 with 10ms (1200 Frames) /Repository/IRTC_Data/default/irtc_00004_cube.fit

  • Results for irtc_00003_cube.fit
    • Center (X,Y) = (163.75,130.5)
    • Center of Rotation of AGw#4 (irtc_00004_stacked.fits):
      irtc_00004_stacked.jpg

Find Guide hotspot for center of WFSC pinhole

Seeing ~0.9"

Humidity changes from 10% to 40% in last half hour

6:30 presetTelescope BS9101 ACTIVE gs=0 Need to reset IE and CA Collimate

Data in /Repository/AGW_Data/091009guider

Guider WFSC Delta X
10 1 0
11 2 +0.5
12 3 +1.0
13 4 +1.1
14 5 +1.2
15 6 0
16 7 -0.5
17 8 -1.0
18 9 -1.1
19 10 -1.3
20 11 -1.5
21 12 -1.7
22 13 -1.8
23 14 -1.9
24 15 -2.1
25 16 -2.3

Guider WFSC Delta Y
26 17 0.0
27 18 +0.5
28 19 +1.0
29 20 +1.5
30 21 +2.0
31 22 +1.7
32 23 +1.6
33 24 +1.5
34 25 0
35 26 -0.5
36 27 -1.0
37 28 -0.7
38 29 -0.5
39 30 -0.2
40 31 0
41 32 +0.2
42 33 +0.4
43 34 +0.6

Check Collimate at low Elevation (near the Moon)

Seeing ~1.6"

7:15 presetTelescope M5_0202 ACQUIRE gs=0 Star saturated. Move probe out. See on IRTC center with IE=-30 CA=39

7:20 presetTelescope M5_0202 ACTIVE gs=1 Catalog file format error

7:25 presetTelescope M5_1133 ACTIVE gs=5 Collimate Elevation=45

Moon

8:00 idl presetTelescope to the Moon. Works!

8:20 Test Doug's TclTk gui for manual guiding. irc OffsetGuiding does not seem to work. Change to irc OffsetPointing. Works fine.

8:40 Setting for Delta_Ra = 2.3, Delta_Dec=0.4 at 0.2 Hz

9:00 Took two cubes for IRTC data, 100 and 3000 Frames. Saving of buffer took 3-4 minutes. moon_00005_cube.fit and moon_00006_cube.fit. Used 6000 microsecond exposure

9:10 To correct for Coordinate calculated for the MMT need IE = -55 CA =110

Check Seeing

9:15 presetTelescope BS_9107 ACQUIRE gs=0 No star

9:17 presetTelescope M5_1133 ACQUIRE gs=0 set IE=-18 CA=32

9:20 presetTelescope BS_9107 ACQUIRE gs=0 No star

9:23 presetTelescope BS_9107 ACQUIRE gs=0 Oops, wanted 9108

9:24 presetTelescope BS_9108 ACQUIRE gs=0 IE=-20 CA=36

9:30 presetTelescope BS_9108 ACTIVE gs=0 Collimate

Seeing 1.3" to 2.0"

Humidity 56%

Wind 5-7 m/s

Pointing Model Data

9:47 Set all collimation to go to M1

9:48 Aaron turn on PCS log at GUI /home/telescope/091009moon file

9:54 pointto ACT0368 TelMode=TRACK gs=0 did not get single image with header info. cube_07

10:00 pointto ACT0409 no star found

time star cube
10:04 ACT0368 08
10:13 WT10_405 09
10:19 ACT0340 10 11
10:22 WT10_291 12
10:29 WT10_172 13
10:33 WT10_246 14
10:39 WT10_091 15 16
10:43 ACT0161 17
10:47 ACT0267 18
10:50 ACT0201 19
10:54 ACT0320 20
10:57 ACT0333 21
10:59 WT10_288 22
11:02 WT10_271 23

Results Made a pointing model with 14 stars with RMS=7.35. Copied to /home/telescope/TCS/Configuration/PCS/irtcdg.20091009.ptmod

       coeff       change     value    sigma

  1     IA       -296.810   -296.81   38.293
  2     IE        -14.238    -14.24    6.884
  3     NPAE      -35.049    -35.05   38.546
  4     CA        +51.262    +51.26   51.417
  5     AW         -1.207     -1.21    2.249
  6     AN        +20.512    +20.51    2.533
  7     TF        +47.832    +47.83    9.475
  8     HACA2      -3.804     -3.80    3.855

Sky RMS =   7.35
Popn SD =  11.22

  • tpoint Graphs:
    irtcdg.20091009.ptmod.jpg

"Fly me to the Moon" F. Sinatra

11:12 pointto M5_0269 Active gs=5 Collilmate IE=-20 CA=40

Data in /Repository/IRTC_Data/091009

cube exp frames
bg_00029_cube.fit 6ms 100
bg_00030_cube.fit 3ms 100
bg_00031_cube.fit 6ms 100
bg_00032_cube.fit 6ms 100
lcross_00024_cube.fit 3ms 100
lcross_00025_cube.fit 3ms 3000
lcross_00026_cube.fit 3ms 3000
lcross_00027_cube.fit 3ms 3000
lcross_00028_cube.fit 6ms 3000
flat_00043_cube.fit 0.5 sec 100
flat_00044_cube_fit 0.1 sec 100

20 Hz Movie of Impact Requires MPEG4 player (eg QuickTime for Mac OS or vlc for CentOS)

Could not see any plume. MMT confirmed that they did not see anything in real time.

Collimate and Check Seeing

12:00 pointto m5_0162 and collimate

Image Motion of IRTC

13:00 pointto M5_0256 TelMode=Track gs=0 PA=0 South on meridian

  • 0,0 cube 34

  • +10 ra RADEC mode move left on irtc image (35)

  • +10 dec RADEC mode move up on irtc image (36)

  • +10 X DETXY mode move left on irtc image (37)

  • +10 Y DETXY mode move down on irtc image (38)

  • direction and magnitude on IRTC cube39 +20 dec in RADEC mode cube40 (10 frames)

  • direction and magnitude on IRTC cube41 +20 dec in RADEC mode cube42 (100 Frames)
    • Using IDL to stack the cubes and the FIND routine I get
Image X Y
orient_00041_cube.fit 178.1 36.1
orient_00042_cube.fit 181.7 231.2
    • angle = atan(opp/hyp) = atan((181.7-178.1)/(231.2-36.1))*180/pi = -1.05386

  • Took several pictures

13:20 pointto M5_0256 TelMode=Track gs=0 PA=90 South on meridian

  • 0, 0 flat45 pa90_00045_cube.fit

  • +10 ra RADEC mode image move up on IRTC pa90_00046_cube.fit

  • Took several pictures

- DougMiller - 29 Sep 2009

I Attachment Action Size Date Who Comment
daq.20091009.gifgif daq.20091009.gif manage 52 K 10 Oct 2009 - 09:33 JohnHill Plot of telescope and mirror temperatures
irtc_00002_stacked.jpgjpg irtc_00002_stacked.jpg manage 39 K 12 Oct 2009 - 08:51 DougMiller Center of Rotation of AGw#4 (irtc_00002_stacked.fits)
irtc_00003_stacked.jpgjpg irtc_00003_stacked.jpg manage 40 K 12 Oct 2009 - 08:51 DougMiller Center of Rotation of AGw#4 (irtc_00003_stacked.fits)
irtc_00004_stacked.jpgjpg irtc_00004_stacked.jpg manage 41 K 12 Oct 2009 - 08:52 DougMiller Center of Rotation of AGw#4 (irtc_00004_stacked.fits)
irtcdg.20091009.ptmod.jpgjpg irtcdg.20091009.ptmod.jpg manage 41 K 12 Oct 2009 - 08:12 DougMiller tpoint Graphs
lcross_00025.mpgmpg lcross_00025.mpg manage 8 MB 11 Oct 2009 - 06:04 DougMiller 20 Hz Movie of Impact
Topic revision: r15 - 12 Oct 2009, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback